Thanks for your help!!
Jill has $3.50 in nickles and dimes.If she has 50 coins, how many of each type of coin does she have? McGraw Hill 6th edition,8.2, # 56.

Answer by stanbon(75887)

Jill has $3.50 in nickles and dimes.If she has 50 coins, how many of each type of coin does she have?
---------
Let number of nickels be "x"; their value is 5x cents
Number of dimes is "50-x"; their value is 10(50-x)=500-10x cents
EQUATION:
value + value = 350 cents
5x + 500-10x=350
-5x=-150
x=30 (number of nickels)
50-x=20 (number of dims)
Cheers,
Stan H.